# Privacy-Preserving Federated Learning with Differentially Private Hyperdimensional Computing

### v2

@@ 6 Experimental Results
In the end, Figure 10 presents the accuracy comparison between the proposed method, FedHDPrivacy, and other FL frameworks. We benchmarked our framework against FedAvg, FedNova, FedProx, FedAdam, and FedSGD. These other frameworks operate by uploading local updates to the central server and downloading the global model to the clients. For this comparison, we used a Convolutional Neural Network (CNN) model, detailed in Table 2, to structure the FL system. These frameworks are vulnerable to adversaries that could reveal training sample information. In a continuous learning scenario over 10 communication rounds, the accuracy achieved by FedAvg, FedNova, FedProx, FedAdam, and FedSGD was $45.89\%$, $48.89\%$, $47.20\%$, $34.66\%$, and $33.71\%$, respectively—all of which are below $50\%$. Notably, both FedAdam and FedSGD scored below $35\%$, indicating their inability to distinguish between the three classes in the experimental task. In contrast, FedHDPrivacy, with a privacy budget $\epsilon$ of 10, achieved an accuracy of $71.93\%$ after 10 rounds, outperforming all benchmarked frameworks.
Table 2: CNN model architecture
Layer Type Input Size Output Size Kernel Size/Stride
1 Conv2D + ReLU (1, 90, 10) (32, 90, 10) 3x3 / 1
2 Conv2D + ReLU (32, 90, 10) (64, 90, 10) 3x3 / 1
3 MaxPooling (64, 90, 10) (64, 45, 5) 2x2 / 2
4 Fully Connected (FC) + ReLU (64 * 45 * 5) 1024 -
5 Fully Connected (FC) + ReLU 1024 1024 -
6 Fully Connected (FC) 1024 8 -
@@
Furthermore, while the initial accuracy of the proposed framework and others ranged between $33\%$ and $38\%$, the accuracy of FedHDPrivacy steadily increased with each communication round. For instance, the accuracy improved to $48.66\%$, $55.70\%$, $65.75\%$, $67.50\%$, and $71.93\%$ after 2, 4, 6, 8, and 10 rounds, respectively, showcasing its enhanced performance in a continuous learning scenario. Additionally, the overall accuracy improvement after 10 rounds for FedHDPrivacy was $37.33\%$, significantly higher than the best improvement seen in the other frameworks, which was less than $12\%$—for example, FedNova’s accuracy increased from $37.86\%$ to $48.89\%$. The trend across rounds illustrates a gradual slope of improvement in performance in a continuous learning setting for FedHDPrivacy.
Moreover, we also evaluated the accuracy of Federated HyperDimensional computing (FedHD), the proposed method without introducing noise to the local and global models. The results demonstrate that the accuracy of FedHDPrivacy and FedHD are similar in each round, indicating that the amount of noise added to the HD models was carefully controlled to maintain a balance between privacy and accuracy. In other words, only the required noise was introduced in each round. The accuracy drop caused by the addition of DP noise was minimal. For instance, after 10 rounds, the accuracy of FedHDPrivacy was less than $5\%$ lower than FedHD, demonstrating that the privacy-preserving noise had only a small impact on performance while securing sensitive information.
@@ 7 Conclusions and Future Work
For future work, exploring additional security measures, such as defenses against model posing and free-riding attacks, could further enhance the reliability of the system. Addressing these issues would help prevent dishonest participation in the FL process and improve the accuracy of the final model. Additionally, as IoT devices often record diverse data types, such as images, sensor signals, and textual data, expanding the FedHDPrivacy framework to handle multimodal data fusion—combining different data types for improved model performance—could offer significant advancements. Another promising avenue for future exploration would be the application of FedHDPrivacy in more complex IoT environments, such as edge computing, where low-latency and real-time processing are critical. Furthermore, enhancing the framework to adapt to different privacy budgets dynamically, based on varying data sensitivity across clients, could provide more tailored privacy solutions. Finally, investigating the impact of FedHDPrivacy on energy consumption and computational efficiency in resource-constrained IoT devices would provide insights into optimizing performance for large-scale deployments.
CRediT authorship contribution statement
Fardin Jalil Piran: Methodology, Software, Formal Analysis, Writing - Original Draft, Zhiling Chen: Writing – Review & Editing, Mohsen Imani: Writing - Review & Editing, Funding Acquisition, Funding Acquisition, Farhad Imani: Conceptualization, Supervision, Writing – Review & Editing, Funding Acquisition. All authors read and approve the final manuscript.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v3

@@ 6.2 Benchmark Comparison
Figure 12 presents the accuracy comparison between the proposed method, FedHDPrivacy, and other FL frameworks. We benchmarked our framework against FedAvg, FedNova, FedProx, and FedOpt. Figure 12(a) illustrates the comparison with FL frameworks integrated with a transformer over communication rounds. In a continuous learning scenario spanning 10 communication rounds, the accuracy achieved by FedAvg, FedNova, FedProx, and FedOpt was $46.64\%$, $44.65\%$, $36.31\%$, and $33.32\%$, respectively—all falling below $50\%$. Notably, both FedProx and FedOpt performed poorly, with accuracy below $37\%$, highlighting their limited ability to distinguish between the three classes in the experimental task. In contrast, FedHDPrivacy, with a privacy budget $\epsilon=10$, achieved an accuracy of $70.73\%$ after 10 rounds, significantly outperforming all benchmarked frameworks.
Figure 13: Training time comparison of FedHDPrivacy and benchmark FL frameworks using a transformer and a convolutional neural network. FedHDPrivacy achieves faster training compared to FedAvg, FedNova, FedProx, and FedOpt.
@@
While the initial accuracy of all frameworks ranged between $33\%$ and $42\%$, the accuracy of FedHDPrivacy steadily increased with each communication round. For example, FedHDPrivacy’s accuracy improved to $52.29\%$, $60.00\%$, $63.04\%$, $67.80\%$, and $70.73\%$ after 2, 4, 6, 8, and 10 rounds, respectively. This steady improvement highlights the superior performance of the proposed framework in continuous learning scenarios. Moreover, FedHDPrivacy demonstrated an overall accuracy improvement of $28.98\%$ after 10 rounds, significantly exceeding the best improvement seen in the other frameworks, which was less than $11\%$. For instance, FedNova’s accuracy only increased from $36.02\%$ to $44.65\%$. This trend underscores FedHDPrivacy’s ability to achieve consistent and notable accuracy gains over communication rounds.
Additionally, Figure 12(b) depicts the comparison with FL frameworks integrated with a CNN over communication rounds. Once again, FedHDPrivacy consistently outperformed all other frameworks. While FedNova exhibited the closest performance to the proposed method, its accuracy remained significantly lower. For example, FedNova’s accuracy was $9.97\%$, $13.77\%$, $13.53\%$, $16.16\%$, and $15.95\%$ lower than FedHDPrivacy after 2, 4, 6, 8, and 10 rounds, respectively. These results demonstrate the superior accuracy of FedHDPrivacy compared to benchmarked frameworks in both transformer-based and CNN-based FL settings.
Besides accuracy, we also evaluate the False Positive Rate (FPR) and False Negative Rate (FNR) for FedHDPrivacy and other benchmarks after the training process, as presented in Table 3. These metrics provide insights into the mistakes made by the models in incorrect predictions. The FPR for FedHDPrivacy is 14.63%, while the best transformer-based benchmark, FedAvg, shows a significantly higher FPR of 26.68%, highlighting the limitations of benchmark frameworks in accurate predictions. Similarly, the best CNN-based FL framework, FedNova, achieves an FPR of 22.61%, which is nearly double that of FedHDPrivacy.
For the FNR, FedHDPrivacy achieves a value of 29.27%, significantly lower than the best CNN-based benchmark, FedNova, which has an FNR of 45.22%. Notably, the worst-performing framework, FedOpt, exhibits an FNR of 66.67%, more than double the FNR of FedHDPrivacy. These results emphasize the superior ability of FedHDPrivacy to minimize incorrect predictions compared to the benchmark frameworks.
Figure 13 compares the training time of FedHDPrivacy with other FL frameworks. The training time for FedHDPrivacy is 45.02 seconds. In contrast, the training time for other FL frameworks is at least 2.36 times that of the proposed method, as seen in the case of the CNN-based FedAvg. The training time for CNN-based models is generally lower than transformer-based models due to the computational overhead of transformers. Notably, the training time for transformer-based models is up to 12.48 times that of FedHDPrivacy, as observed for FedProx, highlighting the efficiency and speed of the proposed method’s training procedure.
Energy efficiency is a critical aspect of FL frameworks, particularly in IoT systems where devices often operate under strict energy constraints. Since the training process primarily occurs on the clients, we evaluate the energy consumption of clients across different FL frameworks while considering both transformer- and CNN-based architectures.
Figure 14(a) highlights the energy consumption of FedHDPrivacy compared to transformer-based FL frameworks. A single client in the FedHDPrivacy framework consumes 5.62 KJ, showcasing its energy-efficient design. In contrast, clients in FedAvg, FedNova, FedProx, and FedOpt consume 4.91, 3.99, 5.37, and 4.82 times more energy, respectively. The high energy efficiency of FedHDPrivacy stems from its computationally lightweight HD model, which reduces the overhead typically associated with transformer-based architectures.
Similarly, Figure 14(b) presents the energy consumption comparison for CNN-based FL frameworks. While CNN-based models are generally more energy-efficient than transformer-based ones, FedHDPrivacy continues to demonstrate superior energy efficiency. Clients in FedAvg, FedNova, FedProx, and FedOpt consume 1.79, 1.82, 2.83, and 1.79 times more energy, respectively, compared to FedHDPrivacy. This consistent performance advantage highlights the framework’s suitability for deployment in resource-constrained environments.
The results across all evaluation metrics—accuracy, training time, and energy efficiency—clearly demonstrate the superiority of FedHDPrivacy over traditional FL frameworks. By effectively balancing privacy, computational efficiency, and model performance, FedHDPrivacy provides a robust solution for FL in dynamic IoT environments.
Figure 14: Client energy consumption comparison of FedHDPrivacy and benchmark FL frameworks using (a) transformer and (b) convolutional neural network. FedHDPrivacy demonstrates superior energy efficiency, consuming less energy compared to FedAvg, FedNova, FedProx, and FedOpt.
@@ 7 Conclusions and Future Work
For future work, exploring additional security measures, such as defenses against model posing and free-riding attacks, could further enhance the reliability of the system. Addressing these issues would help prevent dishonest participation in the FL process and improve the accuracy of the final model. Additionally, as IoT devices often record diverse data types, such as images, sensor signals, and textual data, expanding the FedHDPrivacy framework to handle multimodal data fusion—combining different data types for improved model performance—could offer significant advancements. Another promising avenue for future exploration would be the application of FedHDPrivacy in more complex IoT environments, such as edge computing, where low-latency and real-time processing are critical. Finally, enhancing the framework to adapt to different privacy budgets dynamically, based on varying data sensitivity across clients, could provide more tailored privacy solutions.
@@ Acknowledgment
Supplementary Materials
The code used in this study is openly available for reproducibility and can be accessed at the following link: https://github.com/FardinJalilPiran/FedHDPrivacy.